OpenStudy (anonymous):

What is the molarity of 70.5% nitric acid solution

OpenStudy (nechirwan):

That means 70.5% is nitric acid, the other 29.5% is water. Just convert this to grams: For every 70.5 grams of nitric acid, there are 29.5 grams of water. 70.5g x (1molHNO3/63.02g) = 1.118692479 mol HNO3 29.5g (1mL/1g) x (1L/1000mL) = 0.0295 L Molarity: mol of solute/liter of solution 1.118692479 mol HNO3/0.0295 L = 37.9 M Answer: 37.9 M HNO3 Source: Chemistry Honors ~